The redox reaction is O + 2e- ⇌ R, which the concentrations of O and R are 0.3 mM and 0.2 mM, respectively. The exchange current and the transfer coefficient are 3 mA cm-2 and 0.5, respectively, at 298 K. Calculate the following questions, assuming that diffusion is very fast (6%).
a) k0
b) The exchange current for the concentrations of O and R at 0.4 mM and 0.3 mM, respectively.
1
Expert's answer
2020-06-25T07:45:25-0400

k0 = 0.5 /0.3 x 0.2 = 8,33333335.

0.4 - 0.3 = 0.1